### v1

# On the Robustness of Kolmogorov-Arnold Networks: An Adversarial Perspective

Abstract
Kolmogorov-Arnold Networks (KANs) have recently emerged as a novel approach to function approximation, demonstrating remarkable potential in various domains. Despite their theoretical promise, the robustness of KANs under adversarial conditions has yet to be thoroughly examined. In this paper, we explore the adversarial robustness of KANs, with a particular focus on image classification tasks. We assess the performance of KANs against standard white-box adversarial attacks, comparing their resilience to that of established neural network architectures. Further, we investigate the transferability of adversarial examples between KANs and Multilayer Perceptron (MLPs), deriving critical insights into the unique vulnerabilities of KANs. Our experiments use the MNIST, FashionMNIST, and KMNIST datasets, providing a comprehensive evaluation of KANs in adversarial scenarios. This work offers the first in-depth analysis of security in KANs, laying the groundwork for future research in this emerging field.
1 Introduction
In the rapidly evolving field of deep learning, the robustness of neural networks against adversarial attacks has emerged as a cornerstone of research, driven by the ubiquity of their application, often in sensitive areas. As a result, a multitude of methods have been developed to detect and mitigate adversarial attacks, underscoring the critical need for robust defenses in real-world scenarios. These methods range from adversarial training [31, 4, 39, 2] and defensive distillation [36, 35, 6] to approaches like feature squeezing [52], input transformations [17, 18], and the use of auxiliary models for detection purposes [55, 37, 27]. The proliferation of these techniques reflects the ongoing efforts within the research community to fortify neural networks against increasingly sophisticated attacks.
Multilayer Perceptrons (MLPs)—a foundational model in deep learning—have been extensively studied for their vulnerability and defense mechanisms against adversarial attacks [12, 45, 33, 21, 20]. Recently, the introduction of Kolmogorov-Arnold Networks (KANs) [29, 5, 15, 43, 22] has opened new avenues in function approximation, with theoretical promises of enhanced performance and efficiency.
Many studies have compared the performance of MLPs and KANs across a range of tasks. For example, Yu et al. [53] found that MLPs generally perform better in areas such as machine learning and computer vision. Conversely, KANs excel in representing symbolic formulas.
Another interesting comparison was done by Zeng et al. [54], focusing on the performance of KANs and MLPs in handling irregular or noisy functions. They found that KANs exhibited superior performance both for regular and irregular functions. However, in cases involving functions with jump discontinuities or singularities, MLPs demonstrated greater efficacy.
Recently, Dong et al. [13] studied KANs for time-series classification. They concluded that KANs exhibit significant robustness advantages, attributed to their lower Lipschitz constants. Note that they focused on time-series problems, while we focus herein on image tasks; further, they deployed a single attack algorithm (Projected Gradient Descent, PGD [31]), whereas we will deploy four attack algorithms (including PGD).
This paper delves into the robustness or KANs and MLPs when faced with adversarial attacks under white-box conditions, and investigates the transferability of adversarial examples between these two architectures. By examining how adversarial examples crafted for one model type might affect another, we aim to uncover unique susceptibilities associated with KANs. Using the MNIST, KMNIST, and FashionMNIST datasets, our analysis not only sheds light on the robustness of KANs but also contributes to the broader understanding of neural network security in adversarial settings.
In the next section we provide some background on MLPs, KANs, and robustness. Section 3 studies the robustness of KANs against white-box attacks. Section 4 peruses the question: How transferable are adversarial examples created by KANs? Finally, we end with concluding remarks in Section 5.
2 Background
2.1 Multilayer Perceptrons
Multilayer Perceptrons (MLPs) are a class of feedforward artificial neural networks (ANNs) that consist of multiple layers of nodes, typically organized into an input layer, one or more hidden layers, and an output layer. Each node in one layer is connected to every node in the subsequent layer, with weights assigned with these connections, which are adjusted during the training process. MLPs are capable of approximating complex functions and have been widely used in various fields, such as image recognition, speech analysis, and natural language processing.
Mathematically, let $\mathbf{x}\in\mathbb{R}^{d}$ represent the input vector. Define the output of the $i$-th layer as $\mathbf{h}_{i}$, with $\mathbf{h}_{0}=\mathbf{x}$ for the input layer. Let $\mathbf{W}_{i}\in\mathbb{R}^{n_{i}\times n_{i-1}}$ and $\mathbf{b}_{i}\in\mathbb{R}^{n_{i}}$ be the weight matrix and bias vector for the $i$-th layer, respectively, where $n_{i}$ denotes the number of nodes in layer $i$. The output of the $i$-th layer, $\mathbf{h}_{i}$, can be expressed as:
$\mathbf{h}_{i}=\sigma(\mathbf{W}_{i}\mathbf{h}_{i-1}+\mathbf{b}_{i}),$
where $\sigma(\cdot)$ denotes the activation function, which is often ReLU [34] (Rectified Linear Unit) or sigmoid. The final output of the MLP, denoted as $\mathbf{y}$, is given by the output of the last layer $L$:
$\mathbf{y}=\mathbf{h}_{L}=\sigma(\mathbf{W}_{L}\mathbf{h}_{L-1}+\mathbf{b}_{L}).$
Training an MLP involves optimizing the weights $\{\mathbf{W}_{i}\}$ and biases $\{\mathbf{b}_{i}\}$ to minimize a loss function, typically using gradient-based methods, such as backpropagation.
The universal approximation theorem, established by Cybenko [9] and Hornik et al. [19], states that an MLP with a single hidden layer and a finite number of neurons can approximate any continuous function on a compact subset of $\mathbb{R}^{n}$ with suitable weights and activation functions.
2.2 Kolmogorov-Arnold Networks
Kolmogorov Arnold Networks (KANs) are based on the Kolmogorov-Arnold representation theorem [23], which asserts that any multivariate continuous function can be represented as a finite superposition of univariate functions and additions. KANs leverage this theorem to construct neural networks with a specific architecture that is different from standard MLPs.
Mathematically, given a continuous function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$, the Kolmogorov-Arnold theorem states that there exist continuous functions $\{\varphi_{i}\}$ and $\{\psi_{ij}\}$ such that:
$f(\mathbf{x})=\sum_{i=1}^{2d+1}\varphi_{i}\left(\sum_{j=1}^{d}\psi_{ij}(x_{j})\right).$
In the context of KANs, the $\varphi_{i}$ functions correspond to the activation functions applied at the hidden nodes, and the $\psi_{ij}$ functions represent the preprocessing applied to the inputs. This leads to a network architecture where the input is first transformed by a set of univariate functions, followed by linear combinations and further nonlinear transformations.
KANs offer a different approach to function approximation compared to MLPs, potentially providing advantages in terms of expressiveness and robustness under certain conditions.
KANs leverage the aforementioned theorem to construct neural-network architectures that decompose complex functions into simpler, univariate components. This approach contrasts with the layer-by-layer processing in MLPs, offering potential advantages in terms of interpretability and parameter efficiency.
The structure of KANs suggests that they might be inherently more robust to certain types of perturbations compared to traditional MLPs, as the decomposition into univariate functions can isolate and mitigate the impact of input variations.
2.3 Robustness in Artificial Neural Networks
Robustness in artificial neural networks (ANNs) refers to their ability to maintain performance when subjected to various forms of perturbations, such as noise, adversarial attacks, or distributional shifts. A robust ANN should not only perform well on the training and validation datasets but also generalize effectively to unseen data, including those with slight modifications or corruptions.
Mathematically, robustness can be defined as follows: Consider an ANN $f_{\theta}:\mathbb{R}^{n}\to\mathbb{R}^{m}$, parameterized by $\theta$, and let $\mathbf{x}\in\mathbb{R}^{n}$ be an input to the network. The network is said to be robust to perturbations if, for a small perturbation $\mathbf{\delta}\in\mathbb{R}^{n}$, the output of the network remains unchanged, that is:
$\|f_{\theta}(\mathbf{x})-f_{\theta}(\mathbf{x}+\mathbf{\delta})\|\leq\epsilon\,,$
for some small $\epsilon>0$ and for all $\|\delta\|\leq\eta$, where $\eta$ represents the maximum allowable perturbation norm.
Adversarial robustness in particular has gained significant attention in recent years. Adversarial attacks involve making small—often imperceptible—perturbations to input data that can cause a neural network to produce incorrect outputs with high confidence [16]. These attacks exploit the vulnerability of neural networks to specific input manipulations, raising concerns about their reliability in real-world applications [41].
For example, the Fast Gradient Sign Method (FGSM) demonstrates how slight changes to image pixels can deceive a classifier into misclassification [16]. Additionally, the Carlini-Wagner (C&W) attack shows that even robust defenses can be circumvented through optimized perturbations [7].
Such examples underscore the critical need for developing networks with enhanced robustness against adversarial attacks. Numerous other attack methodologies further highlight this necessity, revealing the extensive landscape of adversarial vulnerability in neural networks [16, 28, 41, 42, 48, 26, 8, 7, 25, 3, 24].
Several techniques have been proposed to enhance the robustness of neural networks, including adversarial training, regularization methods, and architectural modifications. Adversarial training involves training the network on adversarially perturbed examples, thereby improving its ability to withstand attacks [31]. Regularization methods—such as dropout [40] and weight decay—help prevent overfitting and improve generalization. Architectural modifications—such as incorporating attention mechanisms [44] or leveraging alternative network structures—also contribute to robustness.
The upshot is that robustness is a critical aspect of ANN performance, particularly in the context of adversarial attacks and other perturbations. By understanding and improving the robustness of ANNs, we can develop more reliable and secure models for various applications.
3 White-Box Attacks on KANs
Do KANs offer enhanced security compared to MLPs when confronted with a white-box adversary?
To address this question we conduct an experimental evaluation of KANs. While it may appear unconventional to begin with experiments, direct assessment of the security of KANs is most effectively achieved through targeted attacks and subsequent analyses. We begin our study of KAN robustness with a white-box adversary because it represents the most formidable threat model.
3.1 Threat Model
We consider an adversary with full knowledge of the model’s architecture and trained parameters. We assume that the adversary can perturb the original input $\mathbf{x}$ to generate an adversarial example $\mathbf{x_{\text{adv}}}$ within a specified perturbation bound, such that $\|\mathbf{x}-\mathbf{x_{\text{adv}}}\|_{\infty}\leq\epsilon$. For all the datasets, we set $\epsilon=32/255$.
The adversary is considered successful if it can generate an adversarial input $\mathbf{x_{\text{adv}}}$ within this bound that the classifier misclassifies. This scenario is known as an untargeted attack.
Formally, given a classifier $f_{\theta}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ and an original input $\mathbf{x}$ with ground truth label $y$, the goal of the untargeted adversary is to find an adversarial example $\mathbf{x_{\text{adv}}}$ such that:
maximize $\mathcal{L}(f(\mathbf{x_{\text{adv}}}),y)$
subject to $\|\mathbf{x_{\text{adv}}}-\mathbf{x}\|_{\infty}\leq\epsilon,$
where $\mathcal{L}(\cdot)$ is the loss function used to train the classifier. The adversarial example $\mathbf{x_{\text{adv}}}$ is successful if the classifier’s prediction on $\mathbf{x_{\text{adv}}}$ differs from the true label $y$, i.e., $f(\mathbf{x_{\text{adv}}})\neq y$. This formulation encapsulates the adversary’s objective of inducing a misclassification while adhering to the perturbation constraints.
To evaluate a model’s robustness, we first select a set of clean test examples that are correctly classified by the model. We then generate adversarial examples from this set using one of four different attacks. The security of the model is quantified by the percentage of examples that remain correctly classified under these adversarial conditions. Given the relatively recent introduction of KANs, we explore a broad range of attacks and models in our experiments. Below, we enumerate the attacks and models utilized in this study, along with the rationale for their inclusion.
3.2 White-Box Attacks
We evaluate the robustness of our models against four distinct white-box attacks. We begin with the Fast Gradient Sign Method (FGSM) [16], a foundational attack that serves as an initial test of model resilience. Subsequently, we assess the models using more potent iterative attacks, including the Projected Gradient Descent (PGD) [31], Carlini & Wagner (C&W) [7], and Momentum Iterative Method (MIM) [14]. The hyperparameters used for these attacks are outlined in Table 1. A detailed description of each attack is provided in Appendix B.
Table 1: Hyperparameters employed in experiments with white-box attacks.
Attack Parameters
FGSM $\epsilon=32/255$
PGD $\epsilon=32/255$, steps = 30, $\alpha=0.01$
C&W $\epsilon=32/255$, steps = 30, $\alpha=0.01$, $c=10$
MIM $\epsilon=32/255$, steps = 30, $\alpha=0.01$
3.3 Classifier Models
We used six different neural network models, comprising three KAN architectures and three MLP architectures. These models were categorized into small, medium, and large configurations, as specified in Table 2. The KAN models were configured with a uniform setting of num knots = 5 and spline order = 3 across all sizes. Each model was trained on the MNIST [11], FashionMNIST [50], and KMNIST [38] datasets for 20 epochs. Training was performed using the AdamW [30] optimizer, with learning rate $1\times 10^{-4}$, weight decay $5\times 10^{-4}$, and batch size 64.
Table 2: Model configurations.
Model #Params Hidden Layers
[…]
3.4 Analysis of White-Box Attacks
Table 3 shows the results of our experiments of white-box attacks on KANs and MLPs.
Table 3: White-box attacks on KANs and MLPs across MNIST, FashionMNIST, and KMNIST datasets. The attacks are performed using the $\ell_{\infty}$ norm with $\epsilon=32/255$ for all datasets. Robust accuracy is given for each corresponding attack. The “Acc” column refers to the clean accuracy of the model. Boldface values indicate the most robust model in each experiment.
Dataset Model FGSM PGD C&W MIM Acc
MNIST $\text{KAN}_{\text{small}}$ 6.9 0.0 0.0 0.0 96.0
$\text{MLP}_{\text{small}}$ 1.4 0.0 0.5 0.6 97.9
$\text{KAN}_{\text{medium}}$ 9.5 0.0 0.0 0.0 97.6
$\text{MLP}_{\text{medium}}$ 10.8 0.8 0.9 1.3 98.3
$\text{KAN}_{\text{large}}$ 37.8 3.2 4.1 0.9 94.5
$\text{MLP}_{\text{large}}$ 9.6 4.9 0.2 4.9 97.5
FashionMNIST $\text{KAN}_{\text{small}}$ 4.8 0.0 0.0 0.0 86.9
$\text{MLP}_{\text{small}}$ 0.9 0.2 0.4 0.4 88.2
$\text{KAN}_{\text{medium}}$ 3.8 0.0 0.0 0.0 86.7
$\text{MLP}_{\text{medium}}$ 5.2 0.2 0.1 0.3 89.2
$\text{KAN}_{\text{large}}$ 16.7 0.2 0.3 0.0 83.2
$\text{MLP}_{\text{large}}$ 1.5 0.0 0.1 0.1 88.5
KMNIST $\text{KAN}_{\text{small}}$ 2.8 0.0 0.0 0.0 81.7
$\text{MLP}_{\text{small}}$ 3.1 1.2 1.5 1.5 90.3
$\text{KAN}_{\text{medium}}$ 7.6 0.0 0.0 0.0 87.5
$\text{MLP}_{\text{medium}}$ 11.7 1.0 1.1 1.6 91.1
$\text{KAN}_{\text{large}}$ 25.9 0.9 1.2 0.1 78.9
$\text{MLP}_{\text{large}}$ 7.1 0.8 0.3 1.2 89.3
In the context of the iterative adversarial attack methods (PGD, C&W, MIM), small and medium KAN models exhibited lower robustness, when compared with small and medium MLP models as shown in Table 3. Across all examined datasets, medium KAN model displayed a lack of resistance against the deployed attacks. A similar trend was observed in small KAN models, which generally showed almost no robustness, except in the specific case of the MNIST dataset.
In contrast, small and medium MLPs exhibited a certain degree of resistance against these attacks. The sole exception to this trend was observed for the small-sized MLPs with the MNIST dataset, which were completely vulnerable, failing to correctly classify any adversarial images.
The trend whereby MLPs generally exhibit greater robustness compared to KANs changes when considering the larger configurations. For large models, the robustness of KAN improves, surpassing that of large MLP under the C&W attack across all the datasets, and under the PGD attack on FashionMNIST and KMNIST datasets. Not only do large KANs surpass large MLPs in robustness under some iterative-adversarial attacks, but they also exhibited marked improvement when compared to smaller and medium KANs. Across all iterative adversarial attacks and datasets, large KAN showed significantly greater resilience than both small and medium KANs.
Further evidence supporting the robustness of the large KAN model can be observed in Figure 1. The figure illustrates that the loss function (which increases as the discrepancy between the predicted label and the true label is maximized), is consistently lower for the large KAN compared to the other models. This indicates that it is more difficult to ‘deceive’ this model. Additionally, another indicator of this model’s resilience is that, while the losses of other models tend to converge, the loss of the large KAN remains linear.
Figure 1: Comparative loss dynamics of MLPs and KANs under PGD, MIM, and C&W attacks on the MNIST, FashionMNIST, and KMNIST datasets (from top to bottom, respectively). Each line represents the mean loss across batches per iteration, with shaded areas indicating the standard deviation.
In the context of the non-iterative attack method (FGSM), results were somewhat mixed when comparing small and medium models; however, a distinct pattern emerged for larger models, as seen in Table 3: large KAN significantly outperformed large MLP, demonstrating greater robustness to FGSM attacks. To further illustrate the differneces between the medium and large MLPs and KANs, we conducted additional tests using the FGSM attack, focusing on how robustness changes as the $\epsilon$ value increases.
The results, shown in Figure 2, clearly demonstrate that at small $\epsilon$ values, the robustness of both MLPs and KANs appears relatively similar. However, as epsilon increases, the large KAN model is the only one that maintains a significant level of robustness, distinguishing it from medium KAN and both MLP models. Note that at very high $\epsilon$ values medium KAN exhibits greater robustness than even large MLP.
Figure 2: Robust accuracy of $\text{MLP}_{\text{medium}}$, $\text{MLP}_{\text{large}}$, $\text{KAN}_{\text{medium}}$, and $\text{MLP}_{\text{large}}$ as a function of FGSM attack strength, with varying $\epsilon$ values. Each bar represents a different model, and robustness is measured as the accuracy of the model against adversarial examples generated with specific $\epsilon$ values. The x-axis indicates the $\epsilon$ values used for FGSM attacks, while the y-axis shows the corresponding robust accuracy of the models.
Despite the observed improvements in robustness with increased model size, KANs remain notably fragile to adversarial attacks. The results presented in Table 3 and Figure 1 indicate that even the large KAN models, while more resilient than their smaller counterparts and sometimes outperforming large MLPs, still exhibit significant vulnerability under adversarial conditions. This fragility is particularly pronounced in smaller and medium KAN configurations, which show almost no resistance against certain types of attacks, such as the PGD and MIM methods.
Appendix A shows sample image attacks.
4 Transferability Study
How transferable are the adversarial examples created by KANs?
As demonstrated in Section 3, white-box attacks are highly effective at generating adversarial examples that successfully deceive KANs. We now extend the previous analysis by examining the transferability of adversarial examples that are misclassified by KANs. Transferability refers to the phenomenon whereby adversarial examples are misclassified by multiple classifiers.
The transferability of adversarial examples has been extensively documented for various deep-learning architectures. The concept of adversarial-example transferability was first introduced by Szegedy et al. [41]. Subsequent studies have explored the transferability of adversarial examples across different architectures [47, 46, 51, 49, 1, 10, 32]. In this section, we provide a detailed evaluation and analysis of this phenomenon.
4.1 Transferability Metric
Let $\mathcal{A}$ denote the attack method, and let $f_{\theta}$ represent the model under attack. The attack can be formalized as:
$\mathcal{A}_{f_{\theta}}(\mathbf{x},y)=\mathbf{x}_{\text{adv}},$ (1)
where $(\mathbf{x},y)$ corresponds to the input-label pairs.
Let $g_{\phi}$ denote the target model, which is employed to evaluate the effectiveness of the generated adversarial examples. An adversarial example $\textbf{x}_{\text{adv}}$ is considered to have successfully transferred via a given attack if and only if the following conditions hold:
$g_{\phi}(\mathcal{A}_{f_{\theta}}(\textbf{x},y))\neq y\quad\text{and}\quad f_{\theta}(\textbf{x})=g_{\phi}(\textbf{x})=y.$ (2)
To quantify the transferability of adversarial examples between $f_{\theta}$ and $g_{\phi}$ for a specific attack, we define the transferability metric as:
$t_{f_{\theta},g_{\phi}}^{\mathcal{A}}=\frac{1}{m}\sum_{i=1}^{m}\left\{\begin{array}[]{cl}1&\text{if }g_{\phi}(\mathcal{A}_{f_{\theta}}(\mathbf{x}_{i},y_{i}))\neq y_{i}\\ 0&\text{otherwise}\end{array}\right.,$ (3)
where $m$ is the number of samples.
In our experimental evaluation we assessed the transferability of adversarial examples across multiple attack techniques, discussed in the preceding section. For each attack, a corresponding transferability metric was computed, as defined in Equation 3, quantifying the extent to which adversarial examples generated by a particular attack can deceive the evaluator model under varying conditions.
To identify the most effective attack in terms of transferability between the generator model and the evaluator model, we calculated the overall transferability metric, $t^{\text{total}}_{f_{\theta},g_{\phi}}$, as follows:
$t^{\text{total}}_{f_{\theta},g_{\phi}}=\max_{\mathcal{A}\in\{\text{MIM},\text{FGSM},\text{C\&W},\text{PGD}\}}\left(t_{f_{\theta},g_{\phi}}^{\mathcal{A}}\right).$ (4)
This metric quantifies the upper bound of the effectiveness of adversarial attacks in terms of their ability to transfer across models, thereby providing a benchmark for evaluating the robustness of $g_{\phi}$ against the most potent adversarial inputs.
4.2 Transferability Assessment
We focused on quantifying the transferability of adversarial examples across all six model configurations of Table 2. Each model acted both as a generator and an evaluator. To ensure a consistent evaluation basis, we selected a subset of $m=5000$ samples from the validation sets of MNIST, KMNIST, and FashionMNIST, which were correctly classified by both models in the transfer pair: model used for attack, model used for prediction. This subset provided a controlled environment to test the adversarial robustness and transferability of attacks.
The transferability of each pair of models was examined using two equations: Equation 3 was used to assess whether two attacks behaved very differently from each other when applied to the same pair of models; Equation 4 was used to identify the most successful attack and to determine the maximum transferability between two models.
The experimental setup aims to explore transferability not only between different types of models, such as MLPs and KANs, but also among models of varying sizes within the KAN category. This dual focus allows us to discern patterns of vulnerability related to both model architecture and size.
4.3 Analysis of Transfer Attacks
In our analysis of the transferability of adversarial attacks, we observed a consistent pattern: the interaction between the model generating the attack and the model evaluating it remained stable across different types of attacks, as shown in Figure 3 (for KMNIST—below we discuss all three datasets).
Figure 3: Comparison of adversarial attack transferability for the KMNIST dataset, using PGD, C&W, MIM, and FGSM attacks. The x-axis represents the model used for generating the attack, the y-axis shows the model used for prediction, and the z-axis indicates the transferability value. Shades of blue are used to represent scenarios where KANs attack MLPs or vice versa, emphasizing cross-model transferability. Shades of red indicate instances where the type of the attack model is the same as the predictor model, highlighting intra-model transferability.
This observation will further help us understand transferability between KANs and MLPs, and how the size of two different KAN models influence transferability. When adversarial images were generated by an MLP model we observed a markedly higher transferability rate, especially when the evaluator was another MLP model. This suggests that MLPs share common vulnerabilities, making them more susceptible to adversarial examples generated by similar architectures. The homogeneity in their architecture likely contributes to this increased transferability, as adversarial perturbations that exploit weaknesses in one MLP tend to be effective against other MLPs.
In contrast—and somewhat unexpectedly—when adversarial examples were generated by a KAN model, there was no significant increase in transferability between KANs and MLPs. In fact, in some cases, there was more transferability between a KAN-attacked model and an MLP evaluator than between KAN-attacked models and KAN evaluators, as evidenced in Table 4.
Table 4: Maximum transferability for FashionMNIST, KMNIST, and MNIST. The value in row $i$, column $j$ represents the maximum transferability between row-$i$ model and column-$j$ model across four attacks—FGSM, PGD, C&W, MIM—calculated using Equation 4. Unlike robustness tables, the values here indicate the percentage of images misclassified. The ‘Average’ row shows the average transferability for the attacking model, with boldface values denoting the lowest transferability, where a lower value indicates better robustness.
(a) FashionMNIST
$\text{KAN}_{\text{small}}$ $\text{KAN}_{\text{medium}}$ $\text{KAN}_{\text{large}}$ $\text{MLP}_{\text{small}}$ $\text{MLP}_{\text{medium}}$ $\text{MLP}_{\text{large}}$
$\text{KAN}_{\text{small}}$ 100.0 40.5 31.7 81.1 65.8 57.2
$\text{KAN}_{\text{medium}}$ 57.1 100.0 29.2 81.3 69.9 58.3
$\text{KAN}_{\text{large}}$ 38.4 28.1 99.9 59.2 45.1 39.4
$\text{MLP}_{\text{small}}$ 59.5 39.9 31.2 99.7 89.7 75.1
$\text{MLP}_{\text{medium}}$ 55.5 39.9 25.7 98.4 99.8 82.2
$\text{MLP}_{\text{large}}$ 48.2 35.8 26.5 90.8 84.8 100.0
Average ($\downarrow$) 59.78 47.37 40.70 85.08 75.85 68.70
(b) KMNIST
$\text{KAN}_{\text{small}}$ $\text{KAN}_{\text{medium}}$ $\text{KAN}_{\text{large}}$ $\text{MLP}_{\text{small}}$ $\text{MLP}_{\text{medium}}$ $\text{MLP}_{\text{large}}$
$\text{KAN}_{\text{small}}$ 100.0 32.8 18.7 70.7 51.8 45.7
$\text{KAN}_{\text{medium}}$ 29.1 100.0 11.7 62.4 42.9 35.7
$\text{KAN}_{\text{large}}$ 20.7 17.5 99.9 46.4 31.2 31.8
$\text{MLP}_{\text{small}}$ 22.0 16.2 7.4 98.7 83.4 50.2
$\text{MLP}_{\text{medium}}$ 13.2 10.7 4.6 89.8 98.8 43.7
$\text{MLP}_{\text{large}}$ 19.3 17.4 10.8 71.5 59.6 99.6
Average ($\downarrow$) 34.05 32.43 25.52 73.25 61.28 51.12
(c) MNIST
$\text{KAN}_{\text{small}}$ $\text{KAN}_{\text{medium}}$ $\text{KAN}_{\text{large}}$ $\text{MLP}_{\text{small}}$ $\text{MLP}_{\text{medium}}$ $\text{MLP}_{\text{large}}$
$\text{KAN}_{\text{small}}$ 100.0 34.6 21.4 81.4 64.3 58.8
$\text{KAN}_{\text{medium}}$ 27.7 100.0 8.4 68.1 54.6 44.7
$\text{KAN}_{\text{large}}$ 22.2 17.2 99.4 49.0 34.8 31.8
$\text{MLP}_{\text{small}}$ 54.8 37.4 17.8 99.7 95.4 78.4
$\text{MLP}_{\text{medium}}$ 30.1 21.6 7.7 95.8 99.6 72.3
$\text{MLP}_{\text{large}}$ 31.9 25.6 14.0 88.3 84.4 99.9
Average ($\downarrow$) 44.45 39.4 28.12 80.38 72.18 64.32
In this table we also observe the “Average” row, which illustrates the average transferability between the attacked model and each of the evaluator models. We can clearly see that across all datasets the adversarial examples generated by the three KAN models successfully fooled the rest of the models far more concinvingly.
5 Concluding Remarks
The introduction of KANs opens new avenues in adversarial machine learning, presenting new challenges and opportunities for enhancing model security. In our analysis, we explored several significant aspects of KANs that have not been extensively studied before.
First, we assessed the inherent robustness of individual KANs against white-box adversarial attacks. Our results indicate that while KANs initially exhibit vulnerability levels similar to those of traditional MLPs, they demonstrate marked improvement in defense as network size increases. This suggests that KANs could be particularly effective in scenarios where larger models are feasible, and robustness is critical.
Secondly, we investigated the transferability of adversarial examples between KANs and MLPs across six different models on MNIST, FashionMNIST, and KMNIST, revealing that cross-model transferability is generally low, which underscores the unique structural and functional characteristics of KANs compared to MLPs. This low transferability suggests that KANs may process information in fundamentally different ways than MLPs. Our comprehensive experiments and analysis underscore the potential of KANs to enhance security in adversarial machine learning.
However, despite these promising findings, it is important to acknowledge that KANs remain fragile in the face of adversarial attacks. Our analysis shows that even as network size increases, KANs are not impervious to adversarial perturbations. This fragility underscores the necessity for continued research to develop more effective strategies for enhancing the robustness of KANs. Understanding and mitigating these vulnerabilities is essential if KANs are to be reliably deployed in security-critical applications.
As this field continues to evolve, further research is needed to fully understand the capabilities and limitations of KANs, paving the way for their practical use when security and reliability are essential.
Appendix A Image Samples of Attacks
Illustration of PGD, C&W, MIM, and FGSM attacks on a single image from the MNIST dataset (Figure 4). We aim to demonstrate how each attack behaves differently for both MLPs and KANs, and how the model sizes influence the behavior of these attacks. Visually, the perturbations observed for MLPs are very different from those found for KANs.
Figure 4: Illustration of PGD, C&W, MIM, and FGSM attacks on a single image from the MNIST dataset, organized into six rows based on model type and size. The first row corresponds to $\text{MLP}_{\text{large}}$, followed by $\text{KAN}_{\text{large}}$, $\text{MLP}_{\text{Medium}}$, $\text{KAN}_{\text{Medium}}$, $\text{MLP}_{\text{Small}}$, and $\text{KAN}_{\text{Small}}$. Each row displays eight columns, where the left four columns show the adversarial images generated by each attack and the right four columns depict the corresponding perturbations.
Appendix B White-Box Attacks
In this appendix we provide a formal mathematical definition of the white-box adversarial model. Additionally, we present comprehensive descriptions of the white-box attacks evaluated in this study.
B.1 Types of White-Box Attacks
In the context of white-box attacks, the adversary generates an adversarial example $x_{\text{adv}}$ from the original input $x$, using a specific technique, denoted as $\mathcal{A}$. The selection of $\mathcal{A}$ plays a critical role in determining the attack’s success rate, which is defined as the percentage of samples that the classifier $C$ misclassifies. The literature presents a wide array of techniques for crafting white-box attacks. In the following sections, we provide a detailed description of each attack method evaluated in this study.
B.1.1 Fast Gradient Sign Method (FGSM)
The Fast Gradient Sign Method (FGSM) [16] generates adversarial examples by adding perturbations in the direction of the gradient of the loss function. Specifically, an adversarial example $x_{\text{adv}}$ is computed as follows:
$x_{\text{adv}}=x+\epsilon\cdot\text{sign}\left(\nabla_{x}\mathcal{L}(x,y;\theta)\right),$ (5)
where $\mathcal{L}$ denotes the loss function of the classifier $C$, $\nabla_{x}\mathcal{L}(x,y;\theta)$ represents the gradient of the loss with respect to the input $x$, and $\epsilon$ is a small perturbation. Notably, in Equation 5, only the sign of the gradient is used, with the magnitude of the perturbation controlled by $\epsilon$. It is crucial to highlight that FGSM is a single-step attack; the adversary computes the gradient by backpropagating through the model once—and directly applies this gradient to perturb the input $x$.
B.1.2 Projected Gradient Descent (PGD)
The Projected Gradient Descent (PGD) method [31] is an iterative extension of the FGSM, designed to generate more robust adversarial examples. The PGD attack iteratively perturbs the input $x$ by repeatedly applying gradient updates and projecting the perturbed image back onto an $\epsilon$-ball around the original input. The adversarial example $x_{\text{adv}}$ after $k$ iterations is computed as:
$x_{\text{adv}}^{k+1}=\Pi_{\mathcal{B}_{\epsilon}(x)}\left(x_{\text{adv}}^{k}+\alpha\cdot\text{sign}\left(\nabla_{x}\mathcal{L}(x_{\text{adv}}^{k},y;\theta)\right)\right),$ (6)
where $\Pi_{\mathcal{B}_{\epsilon}(x)}(\cdot)$ denotes the projection onto the $\epsilon$-ball centered at $x$, $\alpha$ is the step size, and $k$ indicates the iteration number.
Unlike FGSM, which performs a single gradient step, PGD iteratively refines the adversarial perturbation over multiple steps, thereby producing stronger adversarial examples. The projection step ensures that the perturbation remains within the allowable $\epsilon$-ball, maintaining the proximity of the adversarial example to the original input.
B.1.3 Carlini & Wagner (C&W)
The Carlini & Wagner (C&W) attack [7] is a powerful optimization-based adversarial attack designed to generate adversarial examples by minimizing a specially crafted loss function while ensuring that the perturbation remains imperceptible. When employing the $\ell_{\infty}$ norm, the attack seeks to find a minimal perturbation $\delta$ such that the resulting adversarial example $x_{\text{adv}}=x+\delta$ misleads the classifier. The optimization problem is formulated as follows:
$\min_{\delta}\|\delta\|_{\infty}+c\cdot f(x+\delta),$ (7)
subject to:
$x_{\text{adv}}=\text{clip}(x+\delta,0,1),$
where $f(x_{\text{adv}})$ is a loss function that is specifically designed to ensure the misclassification of the adversarial example $x_{\text{adv}}$. The term $\|\delta\|_{\infty}$ represents the maximum perturbation applied to any individual pixel, and $c$ is a constant that balances the trade-off between minimizing the perturbation and maximizing the loss. The clipping function ensures that the perturbed image remains within the valid input space (i.e., pixel values between 0 and 1).
The attack is typically solved through an iterative optimization procedure, where the perturbation $\delta$ is updated in each iteration to reduce the overall objective function.
B.1.4 Momentum Iterative Method (MIM)
The Momentum Iterative Method (MIM) [14] is an enhancement of the basic iterative methods for generating adversarial examples, incorporating momentum to stabilize the gradient updates and avoid poor local maxima. The MIM attack iteratively updates the adversarial example $x_{\text{adv}}$ by accumulating a momentum term, which helps to amplify gradients that consistently point in the same direction across iterations. The adversarial example $x_{\text{adv}}^{k+1}$ after $k$ iterations is computed as:
$g^{k+1}$ $=\mu\cdot g^{k}+\frac{\nabla_{x}\mathcal{L}(x_{\text{adv}}^{k},y;\theta)}{\|\nabla_{x}\mathcal{L}(x_{\text{adv}}^{k},y;\theta)\|_{1}},$ (8)
$x_{\text{adv}}^{k+1}$ $=\Pi_{\mathcal{B}_{\epsilon}(x)}\left(x_{\text{adv}}^{k}+\alpha\cdot\text{sign}\left(g^{k+1}\right)\right),$ (9)
where $g^{k}$ denotes the accumulated gradient at iteration $k$, $\mu$ is the momentum factor, $\alpha$ is the step size, and $\Pi_{\mathcal{B}_{\epsilon}(x)}(\cdot)$ represents the projection onto the $\epsilon$-ball centered at $x$.
By integrating the momentum term, MIM not only improves the convergence of the attack, but also enhances its effectiveness by consistently updating the perturbation in directions that contribute most to the increase in the loss function. This makes MIM particularly potent in generating adversarial examples that can transfer across different models.